# DMAT: An End-to-End Framework for Joint Atmospheric Turbulence Mitigation and Object Detection

### v1

@@ Abstract
Atmospheric Turbulence (AT) degrades the clarity and accuracy of surveillance imagery, posing challenges not only for visualization quality but also for object classification and scene tracking. Deep learning-based methods have been proposed to improve visual quality, but spatio-temporal distortions remain a significant issue. Although deep learning-based object detection performs well under normal conditions, it struggles to operate effectively on sequences distorted by atmospheric turbulence. In this paper, we propose a novel framework that learns to compensate for distorted features while simultaneously improving visualization and object detection. This end-to-end framework leverages and exchanges knowledge of low-level distorted features in the AT mitigator with semantic features extracted in the object detector. Specifically, in the AT mitigator a 3D Mamba-based structure is used to handle the spatio-temporal displacements and blurring caused by turbulence. Features are extracted in a pyramid manner during the mitigation stage and passed to the detector. Optimization is achieved through back-propagation in both the AT mitigator and object detector. Our proposed DMAT outperforms state-of-the-art AT mitigation and object detection systems up to a 15% improvement on datasets corrupted by generated turbulence.
@@ 1 Introduction
To address the challenges posed by Atmospheric Turbulence in video processing, current learning-based approaches typically adopt a range of strategies. Some methods focus on enhancing the visual quality of videos before applying object detection to the restored outputs [25, 15]. Others retrain existing object detectors on synthetic datasets of atmospheric turbulence [37]. Another strategy includes integrating motion maps, which represent displacements caused by spatial distortions, into the network inputs of the object detectors [29]. Alternatively, some modify the detector’s network architectures to better handle turbulence-induced distortions [20, 16]. In contrast to these methods, we introduce a novel approach that aggregates information in order to enhance feature quality, effectively enabling simultaneous video restoration and object detection within a cohesive, hybrid framework.
Existing methods have achieved considerable success in mitigating atmospheric turbulence; however, they often fail to ensure that restored or enhanced imagery is optimally configured for automatic object detection. A key challenge is that techniques effective in reducing wavy or ripple effects may inadvertently remove crucial features necessary for object identification within video sequences. To address this, this paper proposes a solution called DMAT (Detection and Mitigation of Atmospheric Turbulence), that not only enhances visual quality but also improves object detection accuracy, as demonstrated in Fig. 1, where our combined architecture outperforms individual AT-mitigation and object detection methods. Our approach aims to strike a balance between video enhancement and practical applicability in surveillance systems, ensuring improvement in both domains.
In this paper, we introduce a novel method that leverages 3D deformable convolutions and Structured State Space sequence (S4) models, known as Mamba. This approach addresses spatial displacement across the temporal domain, mitigates distortions such as blur, and enhances contrast. It is specifically designed to restore images affected by atmospheric turbulence and subsequently feed them into an object detector, which outputs object locations and classifications. Our joint framework is trained end-to-end with carefully designed loss functions, enabling the high-level features from the object detector and the low-level features from the turbulence mitigator to mutually enhance each other, resulting in significant improvements in both visual quality and object detection accuracy.
@@
  • A new optimized set of synthetic turbulent videos based on subsets of the COCO dataset.
@@ 2.1 Atmospheric turbulence removal
Atmospheric turbulence is a limiting factor in the imaging of distant objects, degrading visual quality and impairing the performance of automated object recognition and tracking systems. Image sequences captured in turbulent environments often exhibit blurring and rippling effects, accompanied by fluctuations in light intensity across the scene. This phenomenon is particularly pronounced in environments characterized by intense heat, such as deserts, heated road surfaces, and areas surrounding man-made heat sources. Moreover, atmospheric turbulence exacerbates challenges in long-distance imaging applications, such as surveillance, especially when compounded by additional factors such as fog or haze, which further degrade image clarity and quality. Other environmental parameters, including wind speed, altitude, and humidity, also contribute to the severity of these distortions.
Given the spatio-temporal variability of atmospheric turbulence, restoring a scene distorted by atmospheric turbulence is a challenging problem. Traditional methods have involved frame selection, image registration, image fusion, phase alignment and image deblurring [2, 45, 40]. Removing the turbulence distortion from a video containing moving objects is very challenging, as generally multiple frames are used and they are needed to be aligned. Temporal filtering with local weights determined from optical flow is employed to address this by Anantrasirichai et al. [1]. However, artefacts in the transition areas between foreground and background regions can remain. Removing atmospheric turbulence based on single image processing is proposed using ML by Gao et al. [9]. Deep learning techniques to solve this problem are still in their early stages. However, one method reported employs a CNN to support deblurring [28] and another employs multiple frames using a GAN architecture [7]. This however appears only to work well for static scenes.
Similar to de-hazing, physics-inspired models have been widely developed to remove turbulence distortion [17, 18], while complex-valued CNNs have been proposed to exploit phase information [3]. There was also an attempt to use implicit neural representation (INR) to solve this problem, offering tile and blur correction [18]. However, diffusion models outperform on a single image [27], and transformer-based methods remain state-of-the-art for restoring videos [44, 47]. A recent review can be found in [14].
@@ 2.2 Object Detection
Object detection involves both classification and localization, determining the presence of objects and their precise positions within visual inputs. The complexity of object detection arises from challenges such as varying object scales, occlusions, and diverse backgrounds. State-of-the-art object detection methods are most often based on three core components: i) Backbone Network: A deep CNN, often pre-trained on large datasets, serves as the feature extractor. Architectures such as ResNet [13] and MobileNetV4 [30] are commonly employed; ii) Neck: Intermediate layers within the neck component process features from the backbone to enhance semantic information across different scales. Examples include Feature Pyramid Networks (FPN) [21] and Path Aggregation Networks (PANet) [23]; iii) Head/s: The final layers predict bounding boxes and class probabilities. Depending on the model, this can involve anchor-based or anchor-free mechanisms. These models are broadly categorized into two-stage and one-stage detectors.
For two-stage detectors, conventional detectors, exemplified by the Region-based Convolutional Neural Network (R-CNN) family, first generate region proposals and subsequently classify each proposal. The original R-CNN [11] employed a selective search to extract region proposals, followed by feature extraction using a Convolutional Neural Network (CNN) and classification. Enhancements to R-CNN, such as Fast R-CNN [10] and Faster R-CNN [34] improve efficiency by integrating proposal generation and classification into a unified framework.
For one-stage detectors, the detectors process in a single forward pass, making them significantly faster than two-stage detectors. They predict bounding boxes and class probabilities directly, without requiring Region Proposal Networks (RPN). These models are widely used in real-time applications such as autonomous vehicles and surveillance systems. YOLO (You Only Look Once) [31] is one of the most popular detectors that formulate detection as a regression problem. The model divides an image into a grid and predicts bounding boxes and class probabilities in a single forward pass. Variants of YOLO include YOLOv2 [32] introducing anchor boxes, YOLOv3 [33] adding multi-scale detection, YOLOv4 [4] introducing data augmentation and CIoU loss, and recently YOLOv11 [19], incorporating advanced techniques to further boost performance. YOLO is widely used in real-time applications due to its speed and efficiency, though it can struggle with small objects and occlusions.
Transformer-based approaches have also gained attention. The DETR (DEtection TRansformer) [6] eliminates the need for anchor boxes and non-maximum suppression by using bipartite matching. This architecture consists of a transformer encoder-decoder that processes image features extracted from a convolutional backbone. Later, Deformable DETR [46] improves upon the original model by incorporating multi-scale deformable attention, making it more efficient and better suited for detecting small objects and DINO [41]; enhancing convergence speed and accuracy by introducing improved training strategies.
@@ 3 Methodology
Traditionally, models for AT mitigation and object detection are optimized separately. The AT mitigator is typically trained on synthetic datasets due to the absence of ground truth information, where clean videos are pixel-wise registered to the distorted ones. Post-training, the AT mitigator is deployed to generate restored frames that are then utilized to fine-tune the object detector, as illustrated in Fig. 2 (a). In contrast, we propose a joint optimization strategy, as shown in Fig. 2 (b), where high-level features from the object detector guide the AT mitigator in discerning low-level features with semantic significance. Simultaneously, the object detector benefits from receiving more precise features from the AT mitigator, leading to improved overall performance. This synergistic approach promises to significantly advance the efficacy of both video restoration and object detection in atmospherically challenging conditions.
@@ 3.2 AT Mitigation
Figure 4: AT distortion. (a) Frame 1 and 2. (b) Temporal variation showing intensity changes of two pixels over time. Yellow dots visualize the benefit of deformable convolutions.
Registration module. We incorporate a UNet-like architecture with deformable 3D convolutions implemented at all scales in both the encoder and decoder, denoted ‘Deformable Conv3D block’ in Fig. 3. This configuration is specifically designed to effectively estimate pixel shifts at different scales across frames. The deformable 3D convolutions enable pixel mappings associated with the current convolution kernel to extend beyond the conventional grid search area, accommodating spatio-temporal variations. This adaptability ensures that identical objects in different frames, even when subjected to varying turbulent distortions, can be aligned to exhibit consistent features as illustrated in Fig. 4.
The registration module has a depth scale of 4 ($L0$$L3$) with kernel sizes of 3$\times$7$\times$7, 3$\times$7$\times$7, 3$\times$5$\times$5, and 3$\times$3$\times$3. Larger kernels are used in the initial layers to provide a wider field of view. The feature spaces expand to 32, 128, 128, and 256, with 3D batch normalization and max pooling for encoder downsamplers, while trilinear interpolation is used for decoder upsamplers. The output of each scale is a motion field which is then used for frame registration. The registered frames are then fed to the enhancement module.
Enhancement module. The first step consists in feature extraction using 3D convolutions with a kernel size of 3$\times$7$\times$7. The features are then processed in the 3D Mamba-based UNet-like network and simultaneously fed to the object detector, as shown by the red line in Fig.3. We also employ a UNet-like architecture but incorporate 3D Mamba at the encoder, which our studies have shown to outperform the 3D Swin Transformer [24], in agreement with the findings reported in [14]. The Mamba framework is based on the Structured State Space sequence (S4) model. It dynamically adjusts State Space Model (SSM) parameters in response to varying inputs, effectively addressing the common memory and gradient challenges associated with traditional SSM implementations.
@@
where $F_{\text{i}}$ and $F_{\text{o}}$ represent the input and output feature maps of the Res-Mamba block, respectively, within a 3D convolution context. $\text{Conv}_{1}(\cdot)$ is executed with a kernel size of 1$\times$1$\times$1, followed by batch normalization and ReLU activation. SSM($\cdot$) is a selective SSM layer, trained using four augmented feature inputs, each flipped and rotated differently [12]. This sophisticated configuration culminates in the application of two Transformer blocks for optimal channel adjustment and refinement of the final output before passing it to the object detector.
@@ 3.4 Loss function
where $x$ and $y$ are the predicted and true values, and $\epsilon$ is a small constant (e.g., $31e-$ 3) to ensure numerical stability. The Charbonnier loss provides a smooth gradient even for small errors, making it ideally suited for addressing subtle discrepancies in turbulence-impacted images.
@@
$\begin{split}L_{\text{labels}}&=-\frac{1}{N_{pos}+N_{neg}}[\sum^{N_{pos}}t\log p_{\text{ce}}+(1-t)\log(1-p_{\text{ce}})\\ &+\sum^{N_{neg}}p_{\text{ce}}^{\gamma}\log(1-p_{\text{ce}})],\>\>t=p_{\text{ce}}^{\alpha}\text{IoU}^{(1-\alpha)}\end{split}$ (3)
@@
3.5 Alternating training strategy
Several strategies were considered for incorporating these losses within the training process. One approach involved combining the losses into a single, weighted objective function, enabling joint backpropagation through both networks during each training iteration. However, evaluation tests demonstrated that an alternating optimization scheme yielded improved performance. Specifically, dedicating separate training iterations to optimizing the quality of the mitigated output produced more favorable results. In contrast, the use of a combined weighted loss consistently led to degraded mitigation performance, regardless of the chosen weighting parameters. Therefore, we adopt an alternating training strategy for joint optimization of the mitigation and detection modules. Specifically, during odd-numbered iterations, the entire mitigation model is trained while the object detector remains frozen. In even-numbered iterations, we freeze the mitigation model (except for its last 10 layers) and train both the object detector and these unfrozen layers of the mitigator.
@@ 4.1 Synthetic dataset
We employ the COCO2017 dataset [22], extracting various processed subsets to train and evaluate the combined DMAT system. The COCO2017 dataset includes over 330,000 images, where over 200,000 images are labeled with object instances, 80 object categories are annotated with bounding boxes and segmentation masks, and there are over 5 million object instance annotations.
4.1.1 Training, Validation, and Test Sets:
COCO is split into three sets: i) Training Set: Contains 118,000 images and over 2.5 million object instances; ii) Validation Set: Consists of 5,000 images with 80 object categories; iii) Test Set: Composed of 41,000 images. The test set annotations are not publicly available but are used for evaluation. Due to the unavailability of COCO test set annotations we have used the validation dataset for testing in our experiments.
4.1.2 Processed Subsets of the COCO Dataset:
In order to evaluate the developed system for different use-cases, we have subsampled the number of categories into three distinct application subsets:
  • All: All of the 80 categories of the original COCO dataset.
@@ 4.1.2 Processed Subsets of the COCO Dataset:
  • Top10: The top 10 most frequently found categories within the COCO training dataset: Person, Car, Chair, Book, Bottle, Cup, Dining Table, Bowl, Skis, and Handbag.
  • CarPerson: Just the Car and Person classes (intended for surveillance applications).
4.1.3 Reduced Image Resolution:
The combined turbulence and object detection models use significant amounts of GPU memory due to: (i) the processing of spatio-temporal chunks across several hierarchical levels, and (ii) the large amount of memory required for processing images within the visual transformer structure. We have therefore reduced the resolution of all COCO input images in order to fit the processing of the data into the available GPU memory.
To ensure that object bounding boxes are not truncated upon image cropping, we filter all of the training and validation subsets so that only cropped $256\times 256$ regions that wholly contain all objects in the image are included. This significantly reduces the number of images in the training and validation COCO subsets. However, due to the computational complexity of processing videos during training, the resulting number of images remains large but manageable for training. Furthermore, cropping subsets of the dataset has the advantage that all of the images (and therefore videos) being of a consistent resolution ($256\times 256$).
4.1.4 Synthetic AT Distortion:
We utilized the Phase-to-Space (P2S) Transform [26], a turbulence generation system developed at Purdue University, to efficiently simulate imaging through atmospheric turbulence. We evaluated this method as the best performing synthetic turbulence generation method currently available. The P2S method is based on three key steps: i) Reformulating Spatially Varying Convolutions: Converting these convolutions into a set of invariant convolutions with learned basis functions. ii) Learning Basis Functions: Using known turbulence statistical models to derive optimized basis functions. iii) Implementing the P2S Transform: Transforming phase representations into spatial turbulence effects. The output of the P2S transform is labeled as AT (atmospheric transform) in the results.
4.1.5 Final Training and Validation Video Datasets:
After extracting subsets from the COCO dataset, reducing the resolution to $256\times 256$, and generating 50-frame turbulent videos of the same spatial dimensions, we created training and validation video datasets as shown in Table 1 (a similar set of ground truth videos, labeled as noAT in the graphs below, were created that contained 50 identical frames).
Table 1: Number of videos in each sub-sampled COCO dataset.
All Top10 CarPerson
Training 4594 5000 5000
Validation 503 668 831
4.2 Evaluation metrics
4.2.1 Object detection task:
Following the COCO dataset recommended evaluation, we use the standard metric based on the Average Precision (AP) across multiple intersection-over-union (IoU) thresholds. The metric measures the precision and recall of a detection model. The AP is calculated as the area under the Precision-Recall curve. It aggregates the precision-recall values for different IoU thresholds (typically from 0.50 to 0.95 with a step size of 0.05, referred to as AP[0.50-0.95]). This results in a metric that evaluates the model’s performance at various levels of localization precision. We report a Mean Average Precision (mAP). This is a summary metric that combines average precision (AP) across multiple object categories.
4.2.2 AT mitigation task:
We evaluate the performance of AT mitigation using the synthetic dataset; therefore, objective metrics can be used. These include Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) [39], and perceptual similarity metrics like LPIPS [42].

### v3

@@ Abstract
Atmospheric Turbulence (AT) degrades the clarity and accuracy of surveillance imagery, posing challenges not only for visualization quality but also for object classification and scene tracking. Deep learning-based methods have been proposed to improve visual quality, but spatio-temporal distortions remain a significant issue. Although deep learning-based object detection performs well under normal conditions, it struggles to operate effectively on sequences distorted by atmospheric turbulence. In this paper, we propose a novel framework that learns to compensate for distorted features while simultaneously improving visualization and object detection. This end-to-end training strategy leverages and exchanges knowledge of low-level distorted features in the AT mitigator with semantic features extracted in the object detector. Specifically, in the AT mitigator a 3D Mamba-based structure is used to handle the spatio-temporal displacements and blurring caused by turbulence. Optimization is achieved through back-propagation in both the AT mitigator and object detector. Our proposed DMAT outperforms state-of-the-art AT mitigation and object detection systems up to a 15% improvement on datasets corrupted by generated turbulence. The code is available at https://github.com/pui-nantheera/DMAT and datasets are available at https://zenodo.org/records/17673509.
@@ 1 Introduction
To address the challenges posed by atmospheric turbulence in video processing, current learning-based approaches typically adopt a variety of strategies. Some methods focus on enhancing visual quality before applying object detection to the restored outputs [25, 14]. Others retrain existing object detectors on synthetically generated turbulent datasets [33]. JDATT [22] employs an alternative strategy by jointly training the restoration model and object detector, but this can result in suboptimal performance due to competing objectives. Another strategy involves incorporating motion maps—representing displacements caused by spatial distortions—into the input of object detectors [29]. Alternatively, some approaches modify the detector architectures themselves to better handle turbulence-induced distortions [20, 15]. In contrast, we propose a novel approach that aggregates information to enhance feature quality, enabling simultaneous video restoration and object detection (including both classification and localization) within a unified, hybrid framework.
Existing methods have achieved considerable success in mitigating atmospheric turbulence; however, they often fail to ensure that restored or enhanced imagery is optimally configured for automatic object detection. A key challenge is that techniques effective in reducing wavy or ripple effects may inadvertently remove crucial features necessary for object identification within video sequences. To address this, this paper proposes a solution called DMAT (Detection and Mitigation of Atmospheric Turbulence), that not only enhances visual quality but also improves object detection accuracy, as demonstrated in Fig. 1, where our combined architecture outperforms individual AT-mitigation and object detection methods. Our approach aims to strike a balance between video enhancement and applicability in surveillance systems, achieving strong results in both domains.
In this paper, we introduce a novel method that leverages 3D deformable convolutions and Structured State Space sequence (S4) models, known as Mamba. We chose Mamba for its efficiency and ability to model long-range dependencies, and compared to 3D transformers, it yields better PSNR and mAP with fewer parameters as reported in [13]. Our DMAT addresses spatial displacement across the temporal domain, mitigates distortions such as blur, and enhances contrast. It is specifically designed to restore images affected by atmospheric turbulence and subsequently feed them into an object detector, which outputs object locations and classifications. Our joint framework is trained end-to-end with carefully designed loss functions, enabling the high-level features from the object detector and the low-level features from the turbulence mitigator to mutually enhance each other, resulting in significant improvements in both visual quality and object detection accuracy.
@@
  • A new optimized sets of synthetic turbulent videos based on subsets of the COCO and GOT-10k datasets.
@@ 2.1 Atmospheric turbulence removal
Atmospheric turbulence removal is challenging due to its spatio-temporal variability. Traditional methods involve frame selection, registration, fusion, phase alignment, and deblurring [2, 41, 36]. Video-based approaches, like optical flow–guided filtering [1], struggle with artefacts at motion boundaries. ML-based single-image restoration has been explored [9], while early deep learning methods include CNNs for deblurring [28] and GANs for multi-frame restoration [6], though often limited to static scenes. Physics-inspired models [17, 18], complex-valued CNNs [3], and implicit neural representations [18] offer alternatives. Diffusion models show strong results for single images [27], while transformer-based approaches currently lead in video restoration [40, 43, 24] as reported by a recent survey in [13].
@@ 2.2 Object Detection
Object Detection involves classifying and localizing objects within images. Two-stage detectors, such as R-CNN [11], Fast R-CNN [10], and Faster R-CNN [31], first generate region proposals and then classify them, offering high accuracy at the cost of speed. In contrast, one-stage detectors like YOLO [30] perform detection in a single pass, enabling real-time performance. The latest version, YOLOv11 [19], further improves accuracy and speed, though challenges remain with small and occluded objects. Recently, transformer-based approaches such as DETR [5] have emerged, replacing traditional components like anchor boxes and non-maximum suppression with an end-to-end encoder-decoder architecture. Deformable DETR [42] introduces multi-scale attention to improve efficiency and small object detection, while DINO [37] enhances overall accuracy through improved training strategies.
@@ 3 Methodology
Traditionally, models for AT mitigation and object detection are optimized separately. The AT mitigator is typically trained on synthetic datasets due to the absence of ground truth information. Post-training, the AT mitigator is deployed to generate restored frames that are then utilized to fine-tune the object detector, as illustrated in Fig. 2 (a). In contrast, we propose a joint optimization strategy, as shown in Fig. 2 (b), where high-level features from the object detector guide the AT mitigator in discerning low-level features with semantic significance. Simultaneously, the object detector benefits from receiving more precise features from the AT mitigator, leading to improved overall performance. This synergistic approach promises to significantly advance the efficacy of both video restoration and object detection in atmospherically challenging conditions.
@@ 3.2 AT Mitigation
Registration module.
We incorporate a UNet-like architecture with deformable 3D convolutions implemented at all scales in both the encoder and decoder, denoted ‘Deformable Conv3D block’ in Fig. 3. This configuration is specifically designed to effectively estimate pixel shifts at different scales across frames. The deformable 3D convolutions enable pixel mappings associated with the current convolution kernel to extend beyond the conventional grid search area, accommodating spatio-temporal variations. This adaptability ensures that identical objects in different frames, even when subjected to varying turbulent distortions, can be aligned to exhibit consistent features as illustrated in Fig. 4.
[…]
Enhancement module.
[…]
@@
where $F_{\text{i}}$ and $F_{\text{o}}$ represent the input and output feature maps of the Res-Mamba block, respectively, within a 3D convolution context. $\text{Conv}_{1}(\cdot)$ is executed with a kernel size of $\times1$$\times$1, followed by batch normalization and ReLU activation. SSM($\cdot$) is a selective SSM layer, trained using four augmented feature inputs, each flipped and rotated differently [12]. This sophisticated configuration culminates in the application of two Transformer blocks for optimal channel adjustment and refinement of the final output before passing it to the object detector.
Figure 4: AT distortion. (a) Frames 1 and 2. (b) Temporal intensity variation of two pixels. Yellow dots show the benefit of deformable convolutions.
@@ 3.4 Loss function
where $x$ and $y$ are the predicted and true values, and $\epsilon$ is a small constant (e.g., 1e31e-3) to ensure numerical stability. The Charbonnier loss provides a smooth gradient even for small errors, making it ideally suited for addressing subtle discrepancies in turbulence-impacted images.
@@
$\begin{split}L_{\text{labels}}&=-\frac{1}{N_{pos}+N_{neg}}[\sum^{N_{pos}}t\log p_{\text{ce}}\\ &+(1-t)\log(1-p_{\text{ce}})\\ &+\sum^{N_{neg}}p_{\text{ce}}^{\gamma}\log(1-p_{\text{ce}})],\>\>t=p_{\text{ce}}^{\alpha}\text{IoU}^{(1-\alpha)}\end{split}$ (3)
@@ 4.1 Synthetic dataset
As ground truth data for atmospheric turbulence is unavailable, we generate synthetic AT distortions using the Phase-to-Space (P2S) Transform [26], applying them to both static and dynamic datasets designed for object detection. Due to memory limitations, all images and videos are cropped to a resolution of $256\times 256$ pixels.
Mitigator Detector Params (M) Static scenes Dynamic scenes
All classes Top 10 classes Car and person Car and person
All sizes Small All sizes Small All sizes Small All sizes Small
AT YOLO11m 20.1 0.093 0.012 0.040 0.005 0.104 0.022 0.050 0.001
YOLO11x 56.9 0.124 0.029 0.060 0.009 0.148 0.034 0.093 0.007
DETR-Med 41.3 0.092 0.035 0.031 0.003 0.092 0.021 0.043 0.000
DETR-Large 60.0 0.066 0.026 0.035 0.008 0.104 0.025 0.057 0.002
LWDETR-Tiny 12.1 0.158 0.068 0.057 0.000 0.136 0.039 0.062 0.003
LWDETR-Med 28.2 0.192 0.061 0.081 0.017 0.164 0.049 0.069 0.004
DATUM YOLO11m 5.8+20.1 0.307 0.057 0.151 0.033 0.276 0.130 0.122 0.012
YOLO11x 5.8+56.9 0.317 0.060 0.166 0.032 0.291 0.132 0.133 0.013
DETR-Med 5.8+41.3 0.218 0.076 0.108 0.034 0.195 0.074 0.091 0.009
DETR-Large 5.8+60.0 0.243 0.100 0.120 0.051 0.216 0.087 0.102 0.007
LWDETR-Tiny 5.8+12.1 0.293 0.158 0.190 0.074 0.265 0.120 0.113 0.011
LWDETR-Med 5.8+28.2 0.369 0.200 0.231 0.098 0.306 0.140 0.154 0.015
TMT YOLO11m 22.9+20.1 0.276 0.042 0.154 0.028 0.242 0.092 0.103 0.006
YOLO11x 22.9+56.9 0.307 0.038 0.168 0.026 0.256 0.092 0.112 0.008
DETR-Med 22.9+41.3 0.204 0.062 0.101 0.029 0.162 0.045 0.076 0.002
DETR-Large 22.9+60.0 0.224 0.087 0.129 0.040 0.178 0.052 0.093 0.003
LWDETR-Tiny 22.9+12.1 0.297 0.134 0.188 0.074 0.237 0.097 0.101 0.005
LWDETR-Med 22.9+28.2 0.344 0.165 0.226 0.088 0.271 0.119 0.126 0.007
MAMBA YOLO11m 2.8+20.1 0.312 0.046 0.177 0.030 0.285 0.130 0.134 0.013
YOLO11x 2.8+56.9 0.350 0.037 0.191 0.031 0.292 0.128 0.147 0.011
DETR-Med 2.8+41.3 0.227 0.073 0.113 0.031 0.200 0.067 0.098 0.006
DETR-Large 2.8+60.0 0.275 0.097 0.135 0.041 0.217 0.076 0.128 0.008
LWDETR-Tiny 2.8+12.1 0.316 0.164 0.193 0.076 0.270 0.115 0.135 0.012
LWDETR-Med 2.8+28.2 0.364 0.186 0.249 0.105 0.310 0.138 0.167 0.017
DMAT-Tiny (ours) 2.8+12.1 0.374 0.188 0.214 0.108 0.348 0.178 0.214 0.127
DMAT-Small (ours) 2.8+14.6 0.407 0.234 0.276 0.134 0.373 0.194 0.251 0.143
DMAT-Med (ours) 2.8+28.2 0.419 0.234 0.282 0.137 0.385 0.199 0.269 0.152
Table 1: Object detection performance in terms of mAP[0.50-0.95] for AT mitigation methods and detectors across COCO subsets and object sizes. The top three results are highlighted in bold, underline, and italic, respectively. Parameters column shows number of parameters in the object detector and (where applicable) the number of parameters in the AT-mitigator
4.1.1 Static scenes
We use the COCO2017 dataset [21], which includes over 330k images, with 200k labeled for object detection across 80 categories and over 5 million annotated instances. It is divided into training (118k images), validation (5k images), and test (41k images) sets; since test annotations are unavailable, we use the validation set for evaluation. To explore different use cases, we define three application subsets:
i) All: All 80 COCO categories;
@@ 4.1.1 Static scenes
ii) Top10: The 10 most frequent categories—Person, Car, Chair, Book, Bottle, Cup, Dining Table, Bowl, Skis, and Handbag;
iii) CarPerson: Only the Car and Person classes, relevant to surveillance tasks.
For each image, we generate 50-frame turbulent video sequences, with corresponding ground truth videos comprising 50 identical frames. The final datasets include 4,594 training and 503 validation videos for All, 5,000 training and 668 validation videos for Top10, and 5,000 training and 831 validation videos for CarPerson.
4.1.2 Dynamic scenes
We use the GOT-10k dataset [16], which contains over 10k video segments with bounding box annotations under real-world conditions. For our purposes, we select sequences containing Person or Car objects, extract 50-frame clips and apply synthetic distortions to simulate the spatial and temporal characteristics of AT. As the GOT-10k dataset labels only one object per frame, we generate pseudo ground truth using the YOLOv11x detector. In total, 1,761 videos are used for training and 441 for testing.